# Learning Hierarchical Interactive Multi-Object Search for Mobile Manipulation

### v1

@@ V Experimental Evaluations
We extensively evaluate our approach both in simulation and real-world experiments. We aim to answer the following questions:
I) Does the high-level policy learn to take decisions that lead to efficient exploration of the environment, improving over alternative decision rules?
@@ V-A Experimental Setup
We instantiate the task in the iGibson simulator [35]. Each scene contains three cabinets placed randomly across a set of feasible locations. All the doors in the scene are initially in a closed state. We then construct tasks of finding 1-6 target objects, matching the hardest setting in previous work [1, 3]. We randomly place up to three target objects across the free space of the entire apartment and up to three objects inside the cabinets. We use the same eight training scenes as the iGibson challenge and use the remaining seven apartments for evaluation. The embodied agent is a Fetch robot, equipped with a mobile base with a differential drive, a height-adjustable torso, and a 7-DoF arm. We scale the robot’s size by a factor of 0.85 to be able to navigate the narrow corridors of all apartments. Its raw 10-dimensional action space consists of a continuous linear and angular velocity for the base together with the torso and arm-joint velocities. The robot is equipped with an RGB-D camera with a field of view of 79 degrees and a maximum depth of .
@@ V-B Baselines
HIMOS: The hierarchical reinforcement learning approach presented in Section IV.
@@ V-C Simulation Experiments
| w/o IAM |
46.9 |
47.7 |
49.1 |
46.5 |
49.4 |
49.7 |
48.2 |
| HIMOS | 49.1 | 50.5 | 51.4 | 50.9 | 53.5 | 54.2 | 51.6 |
| Model |
1-obj |
2-obj |
3-obj |
4-obj |
5-obj |
6-obj |
Avg 1-6 |
Success |
Greedy |
94.8 |
94.8 |
92.4 |
92.3 |
89.2 |
87.2 |
91.8 |
| SGoLAM+ |
92.5 |
89.6 |
88.5 |
88.6 |
88.7 |
85.8 |
88.6 |
| w/o frontier |
82.2 |
78.2 |
73.8 |
70.8 |
69.7 |
66.5 |
73.5 |
| w/o expl |
@@
43.7 |
46.2 |
47.5 |
48.8 |
50.0 |
46.8 |
| w/o frontier |
32.0 |
38.6 |
42.7 |
43.6 |
44.8 |
45.2 |
41.1 |
| w/o expl |
40.9 |
42.0 |
46.7 |
48.7 |
48.1 |
48.8 |
45.9 |
| w/o IAM |
45.9 |
46.1 |
45.8 |
48.1 |
49.3 |
49.8 |
47.5 |
| HIMOS | 52.2 | 51.5 | 51.8 | 55.7 | 57.4 | 58.4 | 54.5 |
To test the models’ abilities to learn to complete the tasks, we first evaluate them in the seen apartments for variable numbers of target objects. The results are reported in Table I. We find that all the compared models achieve good success rates. All three high-level decision-making variations, greedy, SGoLAM+, HIMOS, are able to make reasonable decisions, demonstrating the benefits of the design of the high-level abstractions discussed in Section IV-C. Furthermore, our proposed method, HIMOS, further improves over the baselines, consistently achieving the highest success rate and the most efficient paths, as measured by the SPL metric.
We then evaluate the models in the unseen apartments. Note that neither the low- nor high-level policies have seen these scenes during training. The results are shown in Table II. We find that the models learned to generalize without any clear generalization gap. The performance is even slightly higher than on the seen apartments, this is in accordance to previous observations [3]. This may be due to the validation split containing potentially simpler scenes than the training scenes. The evaluations on unseen scenes confirm the observations from the training scenes: HIMOS consistently achieves the highest success rate and the best SPL across all the numbers of target objects. Finally, we find that our hierarchical approach scales very well to longer scenarios, with a very small drop in success rates as the number of target objects increases. To find all six objects, the agent often has to explore the majority of the apartments and interact with a large number of articulated objects.
@@ V-D Ablation Study
Exploration Subpolicies: To evaluate the impact of the exploration policies, we compare HIMOS to w/o frontier and w/o expl. We find frontier exploration to have a large impact on success rates. Removing this component reduces the success rate to 73.5%. Removing the local exploration policy leads to a smaller, but nonetheless significant drop of 7.7 ppt in average success rates as well as a clear drop in SPL. This indicates on one hand the different strengths of the two exploration behaviors, and on the other hand that HIMOS learned to use the local exploration policy to increase search efficiency. Again, we find this effect to be consistent across both seen and unseen apartments.
@@
Qualitatively, we find that the agent learned sensible behaviors for the task at hand. Figure 3 depicts example episodes in the unseen scenes. The agent learned to frequently invoke the local exploration policy while the apartment is still largely unexplored, to then use the global exploration policy to navigate to unexplored corners where target objects could be hidden. Generally, areas further away are being used in order to travel faster to certain areas of interest. The high-level policy also frequently uses frontier-based navigation when the exploration policy is stuck in some area. When a target object lies on the way to another relevant navigation point (a frontier, cabinet, or door), the high-level policy learned to navigate directly to the latter, instead of sequentially navigating to the target object and then proceeding. This behavior saves time and improves efficiency.
@@ V-E Real World Experiments
The experiments are performed in an office building covering three rooms connected by a hallway with a total of three doors. The operation space covers roughly square meters. We place three articulated objects, two cabinets with a revolute door, and one drawer in the environment. These objects are never seen during training. We evaluate runs with 1-6 target objects in five different scenarios, for a total of 30 episodes. Each scenario defines new positions for the three articulated objects. We randomly chose which doors start in an open or closed state, and start each episode from the room that the last episode terminated in. We cover glass doors to prevent the agent from directly looking through them. This evaluation tests (i) the generalization abilities of both low- and high-level behaviors to the real world, (ii) the generalization abilities of the local exploration policy and high-level policy to a different robot, and (iii) the high-level policy’s generalization to unseen subpolicies, as we change both the navigation and manipulation modules. This is an important ability for transfers to different robot models and execution requirements. Lastly, (iv) the map representation enables easy transfer to different objects as it only requires a mapping to known semantic and instance colors.
The results of the experiments are summarized in Table III and example episodes are shown in Figure 4 as well as in the supplementary video. The agent successfully completes 76.7% of the episodes, requiring long sequences of autonomous navigation and physical interactions. The high-level policy proves robust to failures in the subpolicies. These include navigation failures if the planner does not find a valid path to a frontier point and manipulation failures in which the mobile manipulation skill failed to grasp the handle of an articulated object. In this case, the agent is capable of re-triggering the interactions after detecting the failure. A few irrecoverable failures occurred that did not allow the agent to continue: reaching a safety limit of the wrist joint during door opening, base collisions, and in two cases repeated failures of the navigation stack from which the agent was unable to recover.
@@ VI Conclusion
We introduced the interactive multi-object search task in which the agent has to manipulate the environment in order to fully explore it, resembling common household settings. We proposed a novel hierarchical reinforcement learning approach capable of solving this complex task in both simulation and the real world. By combining a high-level policy on abstract action spaces with low-level robot behaviors, we are able to perform long-term reasoning while acting in continuous action spaces. Our approach decouples the perception from decision making which allows a seamless transition to unknown and real-world environments on a differing embodiment. In extensive experiments, we demonstrated the capabilities of our approach and the importance of the individual components in ablation studies. In future work, we will investigate the benefits of jointly training the high- and low-behavior and integrate more sophisticated mapping modules to build a semantic map directly from the robot sensors. Further, additional low-level behaviors could extend environment interaction options or perform more goal-oriented active perception actions.

### v3

@@ V Experimental Evaluations
We extensively evaluate our approach both in simulation and real-world experiments. We aim to answer the following:
I) Does the high-level policy learn to make decisions that lead to efficient exploration of the environment, improving over alternative decision rules?
@@ V-A Experimental Setup
We instantiate the task in the iGibson simulator [36], which builds on the PyBullet physics engine. Each scene contains three cabinets placed randomly across a set of feasible locations. All the doors in the scene are initially in a closed state. We then construct tasks of finding 1-6 target objects, matching the hardest setting in previous work [1, 3]. We randomly place up to three target objects across the free space of the entire apartment and up to three objects inside the cabinets. We use the same eight training scenes as the iGibson challenge and use the remaining seven apartments for evaluation. The embodied agent is a Fetch robot, equipped with a mobile base with a differential drive, a height-adjustable torso, and a 7-DoF arm. We scale the robot’s size by a factor of 0.85 to navigate the narrow corridors of all apartments. Its raw 10-dimensional action space consists of a continuous linear and angular velocity for the base together with the torso and arm-joint velocities. The robot is equipped with an RGB-D camera with a field of view of 79 degrees and a maximum depth of .
@@ V-B Baselines
HIMOS: The hierarchical approach presented in Section IV.
@@ V-C Simulation Experiments
60.9 | 55.7 | 52.2 | 56.0 | 57.6 | 63.5 | 65.7 | 68.7 | 60.6 | |
To test the models’ abilities to learn to complete the tasks, we first evaluate them in the seen apartments for variable numbers of target objects. The results are reported in Table I (left). We find that all the compared models achieve good success rates. All three high-level decision-making variations, greedy, SGoLAM+, HIMOS, are able to make reasonable decisions, demonstrating the benefits of the design of the high-level abstractions discussed in Section IV-C. Furthermore, our proposed method, HIMOS, further improves over the baselines, consistently achieving the highest success rate and the most efficient paths, as measured by the SPL metric.
We then evaluate the models in the unseen apartments. Note that neither the low- nor high-level policies have seen these scenes during training. The results are shown in Table I (right). We find that the models learned to generalize without any clear generalization gap. The performance is even slightly higher than on the seen apartments, this is in accordance with previous observations [3]. This may be due to the validation split containing potentially simpler scenes than the training scenes. The evaluations on unseen scenes confirm the observations from the training scenes: HIMOS consistently achieves the highest success rate and the best SPL across all the numbers of target objects. Finally, we find that our hierarchical approach scales very well to longer scenarios, with a very small drop in success rates as the number of target objects increases. To find all six objects, the agent often has to explore the majority of the apartments and interact with a large number of objects.
@@ V-D Ablation Study
Exploration Subpolicies: To evaluate the impact of the exploration policies, we compare HIMOS to w/o frontier and w/o expl. We find frontier exploration to have a large impact on success rates. Removing this component reduces the success rate to 66.5%. Removing the local exploration policy leads to a smaller, but nonetheless significant drop of 12.0 ppt in average success rates as well as a clear drop in SPL. This indicates the different strengths of the two exploration behaviors, as well as that HIMOS learned to use the local exploration policy to increase search efficiency. Again, we find this effect to be consistent across both seen and unseen apartments.
@@
Qualitatively, we find that the agent learned sensible behaviors for the task at hand. Figure 3 depicts example episodes in the unseen scenes. The agent learned to frequently invoke the local exploration policy while the apartment is still largely unexplored, to then use the global exploration policy to navigate to unexplored corners where target objects could be hidden. Generally, areas further away are being used in order to travel faster to certain areas of interest. The high-level policy also frequently uses frontier-based navigation when the exploration policy is stuck in some area. When a target object lies on the way to another relevant navigation point (a frontier, cabinet, or door), the high-level policy learned to navigate directly to the latter, instead of sequentially navigating to the target object and then proceeding. This saves time and improves efficiency.
@@ V-E Real World Experiments
The experiments are performed in an office building covering three rooms connected by a hallway with a total of three doors. The operation space covers roughly square meters. We place three articulated objects, two cabinets with a revolute door, and one drawer in the environment. These objects are never seen during training. We evaluate runs with 1-6 target objects in five different scenarios, for a total of 30 episodes. Each scenario defines new positions for the three articulated objects. We randomly chose which doors start in an open or closed state, and start each episode from the room that the last episode terminated in. We cover glass doors to prevent the agent from directly looking through them. This evaluation tests (i) the generalization abilities of both low- and high-level behaviors to the real world, given accurate semantic perception, (ii) the generalization abilities of the local exploration policy and high-level policy to a different robot, and (iii) the high-level policy’s generalization to unseen subpolicies, as we change both the navigation and manipulation modules. This is an important ability for transfers to different robot models and execution requirements. Lastly, (iv) the map representation enables easy transfer to different objects as it only requires a mapping to known semantic and instance colors. Our aim is to evaluate the transfer of decision making ability within our stated assumptions. We leave full system evaluation with integrated perception pipeline for future work. We provide details on how this could be implemented in the supplementary.
The results of the experiments are summarized in Table II and example episodes are shown in Figure 4 as well as in the supplementary video. The agent successfully completes 76.7% of the episodes, requiring long sequences of autonomous navigation and physical interactions. The high-level policy proves robust to failures in the subpolicies. These include navigation failures if the planner does not find a valid path to a frontier point and manipulation failures in which the mobile manipulation skill fails to grasp the handle of an articulated object. In this case, the agent is capable of re-triggering the interactions after detecting the failure. A few irrecoverable failures occurred: reaching a safety limit of the wrist joint during door opening, base collisions, and in two cases repeated failures of the navigation stack.
@@ VI Conclusion
We introduced the interactive multi-object search task in which the agent has to manipulate the environment in order to fully explore it, resembling common household settings. We proposed a novel hierarchical, reinforcement learning-based approach capable of solving this complex task in both simulation and the real world. By combining a high-level policy on abstract action spaces with low-level robot behaviors, we are able to perform long-term reasoning while acting in continuous action spaces. Our approach decouples the perception from decision making which allows a seamless transition to unknown and real-world environments on a differing embodiment. Tighter integration with perception and active perception in the lower level are promising areas for future work. In extensive experiments, we demonstrated the capabilities of our approach and the importance of the individual components in ablation studies. In future work, we will investigate the benefits of jointly training the high- and low-behavior and integrate more sophisticated mapping modules directly based on the robot sensors. Further, additional low-level behaviors could extend environment interaction options or perform more goal-oriented active perception actions. Detection of more finegrained failure-feedback could further help the high-level decision making. Lastly, the current instance navigation assumes a pre-specified maximum number of possible instances. Regression-based instance selection is a promising avenue to further scale this approach.